What is the value of x in the equation: 4x^2+20x+25

Answer by ewatrrr(24785)

F First terms
O Outside terms +I Inside terms
L Last terms
....
4x^2+20x+25 = 0
(2x + 5)(2x + 5)= 0
x = -5/2